OpenStudy (anonymous):

Is the set V(2x2) matrices of the form [a b / 0 c] spaces with the indicated operations of M22? If not, why not? Thank you in advance :)

OpenStudy (anonymous):

yes, because it satisfies the addition and multiplication axiom if you take \(V_1=\left[\begin {matrix}a&b\\0&c\end{matrix}\right]\) and \(V_2=\left[\begin {matrix}e&f\\0&g\end{matrix}\right]\) you have V1+V2 = V3 in V ( has the same form with 0 at entry 22) if you have a scalar \(\alpha\) and times it to V1, you have the matrix has the same format. That proves what you need

OpenStudy (anonymous):

oh, so basically the V(2x2) in the question is the general form of the matrices? thank you very much, english is not my native language so I didnt really understand the question :D giving a medal ^_^

OpenStudy (anonymous):

English is not my first language either, hahaha... I always misunderstand what others say.

OpenStudy (anonymous):

do you mind doing 1 more plz? :D

OpenStudy (anonymous):

I will mind if it is above my head, hihihi, just post in the new post, friend.

OpenStudy (anonymous):

okie sir, first time being here :)

OpenStudy (anonymous):

close this post, then you can post the new one